Can you please help me with this one?
Solve by substitution. Determine if the equation is dependent, independent or inconsistant.
y = -3x + 19
y = 2x -1
I'm not sure if I'm doing this right.

Answer by Mathtut(3670)

y = -3x + 19.........eq 1
y = 2x -1............eq 2
:
take the value of y from eq 2 and plug it into eq 1
:
2x-1=-3x+19
:
5x=20
:
x=4
:
now take x's found value and plug it into either equationa and solve for y. I will use eq 2
:
y=2(4)-1=7
:
(4,7) is the solution meaning this is a consistent and independent system since the is one a only one solution